Thomas’ Calculus Early Transcendentals 12ed

Homework Assignment for Chapter 07

1. Section 7.2: Problems 13, 19, 21.

2. Section 7.3: Problems 1, 11, 15, 19, 37, 45, 57.

3. Section 7.3: Derive Table 7.8 using formula for derivatives of inverse functions.

4. Section 7.4: Problems 7, 8, 9, 10, 23(a), 24.